Grade upto college level Physical Chemistry

The pressure exerted by 12 g of an ideal gas at temperature toC in a vessel of volume Vlitre is one atm. When the temperature is increased by 10 degrees at the same volume, the pressure increases by 10%. Calculate the temperature t and volume V. (Molecular weight of the gas = 120.)

Shane Macguire
12 Years agoGrade upto college level

2 Answers

Deepak Patra
12 Years ago
(I) Given P = 1 atm, w = 12 g; T = (t + 273)K; V = V litre
(II) If T = t + 10 + 273 = t + 283 K; V = V litre,
P = 1 + 10/100 = 1.1 atm
Using gas equation, PV = w/m RT
Case I. 1 * V = 12/m R (t + 273) ….(1)
Case II. 1.1 * V = 12/m R (t + 283) ….(2)
From (1) and (2), t = - 137oC or t = 100 K
Also from (1), on substituting t and m (120), V = 0.82 litre
Tapajeet Roy
6 Years ago
PV=nRT=> 1 * V=0.1*0.08*(t+273)=>V=0.008(t+273)Again,PV=nRT1.1*0.008(t+273)=0.1*0.08*(t+283)solving we get, t=-173 Now, V=0.008(-173+273) =0.8